Solve the equation (a) z 4 = 24i − 7, (b) z 2 + z −2 = i, (c) z 6 = (1 + z) 6 


1
Expert's answer
2021-08-03T10:16:07-0400

Solution:

(a):

"z^4=24i-7"

Put "z=x+iy"

"\Rightarrow (x+iy)^4=24i-7
\\ \Rightarrow [(x+iy)^2]^2=24i-7
\\ \Rightarrow [x^2+i^2y^2+2ixy]^2=24i-7
\\ \Rightarrow [(x^2-y^2)+2ixy]^2=24i-7
\\ \Rightarrow [(x^2-y^2)^2+(2ixy)^2+2(x^2-y^2)(2ixy)]=24i-7
\\ \Rightarrow (x^4+y^4-2x^2y^2)-4x^2y^2+4ix^3y-4ixy^3=24i-7
\\ \Rightarrow x^4+y^4-6x^2y^2+i(4x^3y-4xy^3)=24i-7"

On comparing both sides,

"x^4+y^4-6x^2y^2=-7,\ 4x^3y-4xy^3=24
\\\Rightarrow (x^2-y^2)^2-(2xy)^2=-7,\ 4xy(x^2-y^2)=24
\\\Rightarrow (x^2-y^2)^2-(2xy)^2=-7\ ...(i),\ (x^2-y^2)=\dfrac{6}{xy}\ ...(ii)"

Put (ii) in (i),

"\\\Rightarrow (\dfrac6{xy})^2-(2xy)^2=-7
\\\Rightarrow \dfrac{36}{x^2y^2}-4x^2y^2=-7"

Put "x^2y^2=t"

"\\\Rightarrow \dfrac{36}{t}-4t=-7"

"\Rightarrow t=-\dfrac{9}{4},\:t=4
\\ \Rightarrow x^2y^2=-\dfrac{9}{4},\:x^2y^2=4"

Rejecting negative value as it is impossible for real numbers.

"\Rightarrow xy=\pm2
\\ \Rightarrow x=\dfrac{\pm2}{y} \ ...(iii)"

Put (iii) in (ii)

"(\dfrac{4}{y^2}-y^2)=\dfrac{6}{\pm2}=\pm3
\\\Rightarrow \dfrac{4}{y^2}-y^2=3; \dfrac{4}{y^2}-y^2=-3
\\ \Rightarrow y=-4,1;y=-1,4
\\ \Rightarrow y=\pm4,\pm1"

Put this in (iii),

"x=\pm\dfrac{1}{2},\pm2"

Thus, solution is "z=\pm\dfrac12\pm4i;z=\pm2\pm1i"

(b):

"z^ 2 + z \u22122 = i"

Put "z=x+iy"

"\Rightarrow (x+iy)^ 2 + (x+iy) −2 = i
\\ \Rightarrow x^2-y^2+2ixy+x+iy-2=i
\\ \Rightarrow (x^2-y^2+x-2)+i(2xy)=0+i"

On comparing,

"x^2-y^2+x-2=0;\ 2xy=1
\\\Rightarrow x^2-y^2+x-2=0\ ...(i);\ y=\dfrac{1}{2x} \ ...(ii)"

Put (ii) in (i),

"\\ \Rightarrow x^2-(\dfrac{1}{2x})^2+x-2=0
\\ \Rightarrow 4x^4-1+4x^3-8x^2=0
\\ \Rightarrow 4x^4+4x^3-8x^2-1=0
\\ \Rightarrow x\approx \:1.07 ,\:x\approx \:-2.02"

Put these in (ii)

"y=0.47,y=-0.25"

Thus, solutions are "z=1.07+0.47i;\ z=-2.02-0.25i"

(c):

"z ^6 = (1 + z)^ 6 
\\\Rightarrow (\dfrac{z}{1+z})^6=1 
\\\Rightarrow \dfrac{z}{1+z}=1
\\\Rightarrow \dfrac{z}{1+z}=e^{ik\pi/3}" , where "k=1,2,..,6"

"\\\Rightarrow z=\dfrac{e^{ik\pi/3}}{1-e^{ik\pi/3}}"

 Simplify this using trigonometric form and double angle identites to get:

"z=-\dfrac12+i\cot(\dfrac{k\pi}{6})"

Putting "k=1,2,...,6"

"z=-\dfrac12+i\cot(\dfrac{\pi}{6});-\dfrac12+i\cot(\dfrac{\pi}{3});-\dfrac12+i\cot(\dfrac{\pi}{2})
\\;-\dfrac12+i\cot(\dfrac{2\pi}{3});-\dfrac12+i\cot(\dfrac{5\pi}{6});-\dfrac12+i\cot(\pi)"

"z=-\dfrac12+i\sqrt3;-\dfrac12+i(\dfrac{1}{\sqrt3});-\dfrac12+i(0)
\\;-\dfrac12+i(-\dfrac{1}{\sqrt3});-\dfrac12+i(-\sqrt3);-\dfrac12+i(\infty)"

Thus, solutions are:

"z=-\dfrac12+i\sqrt3;-\dfrac12+i(\dfrac{1}{\sqrt3});-\dfrac12
\\;-\dfrac12-i(\dfrac{1}{\sqrt3});-\dfrac12-i(\sqrt3);\text{not defined}"